OpenStudy (anonymous):

(a) Find f(x) for the indicated values of x, if possible. (b) Find the domain of f. f(x) = 1/(x^2 + 4) for x = -2, a + 4

OpenStudy (anonymous):

\[f(x) = \frac{1}{x ^{2}+4} for\ x = -2, a + 4\]

OpenStudy (anonymous):

At, x=-2 f(-2)=1/((-2)^2+4) =1/(4+4) =1/8 f(a+4)=1/((a+4)^2+4) =1/(a^2+8a+16+4) =1/(a^2+8a+20) As, for the domain, when we take alook a the denominator of the function we see an x^2 term, what this indicates is that we can plug in any x value and get out a valid y value. Meaning that the domain of this function is all real numbers .